OpenStudy (anonymous):

what is the derivative of log(base 10)x?

OpenStudy (anonymous):

You can write it in terms of the natural logarithm by the change of base formula: \[\log_b a = \frac{\ln a}{\ln b}\]

OpenStudy (anonymous):

Got it! Thanks :)

OpenStudy (kropot72):

The derivation is lengthy, however I will reproduce it if you need it. Let log(base 10)x = y \[\frac{dy}{dx}=\frac{1}{x}\log_{10} e\]

OpenStudy (anonymous):

Surely the derivation isn't long at all..? \[\frac{d}{dx}\log_{10} x = \frac{d}{dx}\frac{\ln x}{\ln 10} = \frac{d}{dx}\frac{\log_{10} e}{x}\]

OpenStudy (anonymous):

Oops no derivative sign on the last one!

OpenStudy (kropot72):

I was referring to a derivation that includes a demonstation that \[\frac{d(\log_{e} x)}{dx}=\frac{1}{x}\]

OpenStudy (anonymous):

Ah okay :)

OpenStudy (agent0smith):

@kropot72 did you mean something like this? http://www.youtube.com/watch?v=yUpDRpkUhf4

OpenStudy (kropot72):

@agent0smith Sort of :)